A ray of light is incident in water at an angle of

(i) 30 degree.

(ii) 70 degree on a water glass plane surface. Calculate the angle of refraction in the glass in each case. ( n of glass = 1.5 , n of water = 1.33)


1
Expert's answer
2021-03-09T15:31:24-0500

(i) Let's write the Snell's Law:


"n_isin\theta_i=n_rsin\theta_r,""\theta_r=sin^{-1}(\dfrac{n_isin\theta_i}{n_r}),""\theta_r=sin^{-1}(\dfrac{1.33\cdot sin30^{\circ}}{1.5})=26.3^{\circ}."


(ii) In case of angle of incidence "70^{\circ}", we get:


"\theta_r=sin^{-1}(\dfrac{1.33\cdot sin70^{\circ}}{1.5})=56.4^{\circ}."